OpenStudy (anonymous):

Determine the number of real solutions of -2x² + 5x - 3 = 0. Answers a.2 b.1 c.0

myininaya (myininaya):

Fine the discriminant \[b^2-4ac\] If <0, then 2 imaginary If >0, then 2 real If =0, then 1 real

OpenStudy (anonymous):

is it A,b,or c,?

myininaya (myininaya):

Plug into the expression above What does a=? b=? c=?

myininaya (myininaya):

Then evaluate the expression

OpenStudy (anonymous):

a,b,c. are the answers

myininaya (myininaya):

I know I'm telling you how to get the answer

myininaya (myininaya):

your equation is in this form \[ax^2+bx+c=0\] So looking at this equation what is a,b, and c?

myininaya (myininaya):

forget about your choices now

myininaya (myininaya):

We are trying to figure out what a, b, and c are in the equation you have

OpenStudy (anonymous):

.3. ok what do i do?

myininaya (myininaya):

Trying to get you to tell me what a, b, and c are in your equation? Your equation is in this form: \[ax^2+bx+c=0\] This is your equation \[-2x^2+5x-3=0\]

myininaya (myininaya):

What is a?

OpenStudy (anonymous):

2

myininaya (myininaya):

no

myininaya (myininaya):

try again

OpenStudy (anonymous):

-2

myininaya (myininaya):

yes! :) now what is b?

OpenStudy (anonymous):

b=5 c=-3

myininaya (myininaya):

Ok great now plug those into the discriminant that I mentioned

myininaya (myininaya):

\[b^2-4ac \]

OpenStudy (anonymous):

5^2-4(-2)(-3)=1

myininaya (myininaya):

OMG great! :) Now I'm going to copy and paste what I said about the discriminant: If <0, then 2 imaginary If >0, then 2 real If =0, then 1 real

myininaya (myininaya):

Which one is it? Is 1 <0? Is 1>0? Is 1=0?

OpenStudy (anonymous):

1=0?

myininaya (myininaya):

No you know 1 is bigger than 0

myininaya (myininaya):

Which on says 1 is bigger than 0?

OpenStudy (anonymous):

oo my bad 1>0

myininaya (myininaya):

Great! :) So The conclusion is ?

myininaya (myininaya):

If >0, then 2 real

OpenStudy (anonymous):

so that answer is A?

myininaya (myininaya):

Yes :)

myininaya (myininaya):

Good job!

OpenStudy (anonymous):

thank you.

myininaya (myininaya):

I will give you a medal for working so hard to figure it out :)

OpenStudy (anonymous):

thanks